# Junyu Cao and Mohsen Bayati

### v1

@@ 7 Concluding Remarks
This work is motivated by the ever-growing observations of human-AI interaction in LLM. We presented a theoretical two-stage framework for explaining the significance of using human comparisons for the model improvement. While the noisy-labeled data is fed into the SFT procedure to learn the low-dimensional representation in the first stage, we sequentially inquire human evaluators to make pair-wise comparisons in the second stage. To address the challenge of efficient acquisition of valuable information from these human comparisons, we introduced a probabilistic bisection method, which factors in the uncertainty arising from the accuracy of human comparisons. The newly introduced concept, label-noise-to-comparison-accuracy ratio, quantifies the relative scale between label noise and human comparison accuracy. We demonstrate the significant advantages of our proposed two-stage framework over pure supervised learning approach under certain conditions.
@@
Our work represents an early attempt to organically integrate machine learning with human knowledge. We consider this paper as a prompt for an open discussion. Some potential extensions to our work are worth investigating. Firstly, we extended the probabilistic bisection algorithm from one dimension to multiple dimensions by refining along each dimension. This expansion raises the question of whether a more efficient bisection algorithm exists in multi-dimensional spaces. As discussed in Frazier et al. (2019), the development of a PBA tailored for multi-dimensional problems remains an open problem. Secondly, although our paper focuses on a prediction problem, the framework has the potential to be extended to the action-based learning, where the goal is to select optimal actions for the decision-making problem. Thirdly, exploring the optimal sample selection to maximize the learning algorithm’s efficiency is a compelling area of inquiry. While we have outlined guidelines for practical sample selection, the quest for the most efficient proposed procedure remains open.

### v2

@@ 8 Discussions and Concluding Remarks
This work is motivated by the ever-growing observations of human-AI interaction in LLM. We presented a theoretical two-stage framework for explaining the significance of using human comparisons for model improvement. In the first stage, the noisy-labeled data is fed into the SL procedure to learn the low-dimensional representation; we then ask human evaluators to make pair-wise comparisons in the second stage. To address the challenge of efficient acquisition of valuable information from these human comparisons, we introduced a probabilistic bisection method, which factors in the uncertainty that arises from the accuracy of human comparisons. The newly introduced concept, the LNCA ratio, quantifies the relative scale between label noise and human comparison accuracy. We demonstrate the significant advantages of our proposed two-stage framework over a pure supervised learning approach under certain conditions.
8.1 Implications for Supervised Fine-Tuning
Supervised fine-tuning (SFT) is the process of refining a pre-trained model on a labeled dataset with specific input-output pairs tailored to a target task. Our proposed framework can fit into SFT in the following way. First, the pre-trained model produces a representation, denoted as $\hat{\varphi}(\cdot)$, along with initial parameters $\hat{\boldsymbol{\theta}}_{0}$. For a new task, if $\hat{\varphi}(\cdot)$ remains the same, we proceed directly to the second phase, where human feedback is solicited for model alignment through comparison data. However, if the representation $\hat{\varphi}(\cdot)$ also requires adaptation, we first update the embeddings by training on a combination of newly collected and existing noisy-labeled data, using techniques such as transfer learning (Pan 2020). In this adapted embedding space, human comparisons are then actively acquired in the second stage to enhance model alignment.
8.2 Future Directions
@@
Our work represents an early attempt to organically integrate two sources of data: noisy-labeled data and human comparison data. We position this paper as a prompt for an open discussion. Some potential extensions to our work are worth investigating. First, we extended the probabilistic bisection algorithm from one dimension to multiple dimensions by refining along each dimension. This expansion raises the question of whether a more efficient bisection algorithm exists in multi-dimensional spaces. As discussed in Frazier et al. (2019), the development of a PBA tailored to multi-dimensional problems remains an open problem. Second, although our paper focuses on a prediction problem, the framework has the potential to be extended to action-based learning, where the goal is to select optimal actions for the decision-making problem. Third, exploring the optimal sample selection to maximize the learning algorithm’s efficiency is a compelling area of inquiry. Although we have offered guidelines for practical sample selection, the quest for the most efficient proposed procedure remains open.